This is an algebraic equation and I am having a hard time finding the answer and doing a check to see if it is right.
3x-7=5(2x-3)+4

Found 2 solutions by John10, solver91311:
Answer by John10(297)

3x-7 = 5(2x-3)+ 4
3x - 7 = 10x - 15 + 4
3x - 7 = 10x - 11
3x - 10x = -11 + 7
-7x = -4
x = 4/7
John10:)

Answer by solver91311(24713)



Distribute the 5 across the parentheses. Collect constant terms in the RHS. Add -10x to both sides. Add 7 to both sides. Multiply both sides by -1/7.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism